# PAS: Data-Efficient Plug-and-Play Prompt Augmentation System

### v2

@@ Abstract.
In recent years, the rise of Large Language Models (LLMs) has spurred a growing demand for plug-and-play AI systems. Among the various AI techniques, prompt engineering stands out as particularly significant. However, users often face challenges in writing prompts due to the steep learning curve and significant time investment, and existing automatic prompt engineering (APE) models can be difficult to use. To address this issue, we propose PAS, an LLM-based plug-and-play APE system. PAS utilizes LLMs trained on high-quality, automatically generated prompt complementary datasets, resulting in exceptional performance. In comprehensive benchmarks, PAS achieves state-of-the-art (SoTA) results compared to previous APE models, with an average improvement of 6.09 points. Moreover, PAS is highly efficient, achieving SoTA performance with only 9000 data points. Additionally, PAS can autonomously generate prompt augmentation data without requiring additional human labor. Its flexibility also allows it to be compatible with all existing LLMs and applicable to a wide range of tasks. PAS excels in human evaluations, underscoring its suitability as a plug-in for users. This combination of high performance, efficiency, and flexibility makes PAS a valuable system for enhancing the usability and effectiveness of LLMs through improved prompt engineering.
@@
Figure 1. We first present the pipeline of the PAS. PAS takes user prompts, enhances them, and then inputs the augmented prompts into LLMs. As illustrated in the second figure, PAS significantly improves responses across all categories in human evaluation.
@@ 1. Introduction
In recent years, rapid advancements in large language models (LLMs)(OpenAI, 2023a; Touvron et al., 2023) have underscored the importance of data management and AI systems in leveraging these technologies(Fernandez et al., 2023; Trummer, 2023; Chen et al., 2023a; Miao et al., 2024; Nie et al., 2023). At the same time, there is an increasing need for scalable plug-and-play systems for LLMs due to their flexibility and efficiency (Al Ridhawi et al., 2020).
Among the various techniques for LLMs, prompt engineering has emerged as a crucial approach due to its extremely low cost and significant enhancement of LLM performance (Sahoo et al., 2024). This technique leverages the inherent capabilities of LLMs to understand and generate human-like text, enabling them to perform a wide range of applications from natural language understanding and generation to specialized domains such as medical diagnosis and legal analysis (Meskó, 2023; Giray, 2023).
In prompt engineering techniques, few-shot learning stands out because it provides a small number of examples to guide the model, thereby enhancing its task-specific performance. Chain-of-Thought (CoT) prompting guides models to reason through problems step-by-step, improving logical consistency and accuracy. In-context learning, by embedding relevant examples and instructions directly within the prompt, allows models to adapt dynamically to new tasks. Despite its potential, prompt engineering is not user-friendly and requires meticulous design. Crafting effective prompts demands a deep understanding of both the model and the task at hand. This process can be time-consuming and often involves extensive trial and error to optimize performance. To tackle those weak points, automatic prompt engineering is designed for easier prompt generation (Zhou et al., 2022a; Pryzant et al., 2023a; Cheng et al., 2023). Although APE models can automatically enhance the prompt, they have to use massive amounts of human-labeled data (Cheng et al., 2023; Ouyang et al., 2022; Rafailov et al., 2024). Additionally, previous methods failed to construct a flexible, user-friendly and effective APE model (Zhou et al., 2022b; Pryzant et al., 2023b). They face the following three key challenges:
@@
  • New Method. We propose a new method that first utilizes diversity and quality selection for prompt construction. Then, few-shot learning is used for prompt complementary dataset construction. We employ GPT to identify and regenerate incorrectly generated data until it is correct. Instead of using human preference data, we only use the prompt complementary dataset to fine-tune LLMs and achieve SoTA performance.
  • SoTA Performance. (1) High Efficiency. Our PAS model requires only 9000 pairs of prompt complementary data to fine-tune LLMs, making it extremely data-efficient. Additionally, our data generation process is entirely automatic and requires no human labor. Moreover, our PAS model can be integrated into any LLM and solve all tasks, achieving SoTA performance. Therefore, we only need to train one LLM, reducing computational costs. (2) High Flexibility. Our PAS model can be plugged into any LLM, and is model and task agnostic. It achieves SoTA performance across all models and tasks, demonstrating its exceptional flexibility. (3) SoTA Performance on Multiple Benchmarks. Our PAS model achieves SoTA performance across multiple models and comprehensive benchmarks. It also outperforms the previous SoTA model, BPO, under the same experimental settings. Furthermore, it achieves superior performance on human evaluation metrics, highlighting the outstanding capabilities and potential uses of the PAS model.
@@ 2.2. Plug-and-Play Systems
In image processing research, plug-and-play systems are commonly applied for its outstanding flexibility. Image reconstruction, denoising, deblurring, image enhancement, and super-resolution are all fields where plug-and-play systems are highly needed. By integrating various image processing modules into a unified framework, plug-and-play systems can flexibly combine different methods to achieve better image processing results. Moreover, this system allows for the easy addition or replacement of new processing modules without the need to redesign the entire algorithm, significantly improving the efficiency and effectiveness of image processing.
@@ 2.3. Data Quality and Data Selection
: High-quality data can significantly enhance the performance of models. (meta llama, 2024) As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation. (Bai et al., 2024) Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
@@
: LLMs-based methods were commonly used in data selection. (Bai et al., 2024) For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023b) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
@@ 3.1. Prompts Data Collection
3.2. Prompts Augmentation Dataset
@@ 4. Experiments
In this section, we first introduce the experimental setups. We then aim to answer the following questions to verify the effectiveness, efficiency, and robustness of our proposed KeyframeLLM: Q1: Can our PAS achieve SoTA performance compared to previous SoTA methods? Q2: Can our PAS outperform the previous SoTA model with the same base model? Q3: How efficient and flexible is our model compared to previous APE models? Q4: Can PAS achieve SoTA performance in human evaluation, making it a user friendly system? Q5: Can we visualize the advantages of our method? Q6: Do we need data quality selection and regenerate module in our data generation pipeline?
@@ 4.1. Experiments Setting
  • Alpaca-Eval 2.0 (LC): This is a variation of the Alpaca-Eval 2.0 benchmark, specifically focusing on language comprehension (LC). It includes tasks that test the model’s ability to understand and interpret complex language inputs accurately. This benchmark is crucial for evaluating the depth of language understanding and the precision of responses generated by the model.
@@ 4.4. Efficiency and Flexibility of PAS
To address Question 3 (Q3), we compare the data efficiency and flexibility of the PAS approach with SoTA APE models. This comparison underscores PAS’s effectiveness in terms of both data utilization and adaptability.
@@ 4.4.1. Data Usage Comparison
Figure 7 illustrates the data consumption of PAS compared to other methods, including PPO, DPO, and BPO. The data consumption for each method is as follows:
  • PAS: 9,000k
  • BPO: 14,000k
  • PPO: 77,000k
  • DPO: 170,000k
As shown in the figure, PAS significantly reduces data consumption compared to other methods, demonstrating its efficiency in data usage. It is important to note that ProTeGi and OPRO are not task-agnostic, so they are not included in the data efficiency comparison.
@@ 4.5. Human Evaluation
To address Q4, we conducted a comprehensive evaluation using human evaluators to assess the online performance of our PAS compared to the baseline model without any prompt augmentation. We compared the grade score benchmark (GSB), availability proportion, full mark proportion, and average score.
@@ 4.6. Case Study
Our PAS approach significantly improves this prompt by providing a complementary hint to remind the model that there is a logic trap. With the assistance of PAS, the new response not only avoids the trap but also showcases a multi-step logical reasoning process. This refined answer is clear, accurate, and guides the user through the reasoning process, explaining why only one bird would be on the ground.
@@
Case Study 3
To summarize, the benefits of our PAS method are as follows:
  • Comprehensive and Clear Responses: The method promotes detailed explanations, ensuring that the response is not only correct but also easy to understand and logically sound, as illustrated in Case Study 3 where PAS provided a thorough analysis of blood pressure changes.

### v5

@@ Abstract.
In recent years, the rise of Large Language Models (LLMs) has spurred a growing demand for plug-and-play AI systems. Among the various AI techniques, prompt engineering stands out as particularly significant. However, users often face challenges in writing prompts due to the steep learning curve and significant time investment, and existing automatic prompt engineering (APE) models can be difficult to use. To address this issue, we propose PAS, an LLM-based plug-and-play APE system. PAS utilizes LLMs trained on high-quality, automatically generated prompt complementary datasets, resulting in exceptional performance. In comprehensive benchmarks, PAS achieves state-of-the-art (SoTA) results compared to previous APE models, with an average improvement of 6.09 points. Moreover, PAS is highly efficient, achieving SoTA performance with only 9000 data points. Additionally, PAS can autonomously generate prompt augmentation data without requiring additional human labor. Its flexibility also allows it to be compatible with all existing LLMs and applicable to a wide range of tasks. PAS excels in human evaluations, underscoring its suitability as a plug-in for users. This combination of high performance, efficiency, and flexibility makes PAS a valuable system for enhancing the usability and effectiveness of LLMs through improved prompt engineering. The code is available at https://github.com/PKU-Baichuan-MLSystemLab/PAS.
@@
Figure 1. We first present the pipeline of the PAS in (a). PAS takes user prompts, enhances them, and then inputs the augmented prompts into LLMs. As illustrated in (b), PAS significantly improves responses across all categories in human evaluation.
@@ 1. Introduction
In recent years, data management and AI systems (Fernandez et al., 2023; Trummer, 2023; Chen et al., 2023a; Miao et al., 2024; Nie et al., 2023) have achieved rapid advancements and play an important role in large language models (LLMs) (OpenAI, 2023a; Touvron et al., 2023). At the same time, there is an increasing need for scalable plug-and-play systems for LLMs due to their flexibility and efficiency (Al Ridhawi et al., 2020).
Among the various techniques for LLMs, prompt engineering has emerged as a crucial approach due to its extremely low cost and significant enhancement of LLM performance (Sahoo et al., 2024). This technique leverages the inherent capabilities of LLMs to understand and generate human-like text, enabling them to perform a wide range of applications from natural language understanding and generation to specialized domains such as medical diagnosis and legal analysis (Meskó, 2023; Giray, 2023). In prompt engineering techniques, few-shot learning (Brown et al., 2020) stands out because it provides a small number of examples to guide the model, thereby enhancing its task-specific performance. Chain-of-Thought (CoT) (Wei et al., 2022) prompting guides models to reason through problems step-by-step, improving logical consistency and accuracy. In-context learning (Dong et al., 2022), by embedding relevant examples and instructions directly within the prompt, allows models to adapt dynamically to new tasks.
Despite the potential of existing methods, prompt engineering is not user-friendly and requires meticulous design. Crafting effective prompts demands a deep understanding of both the model and the task at hand. This process can be time-consuming and often involves extensive trial and error to optimize performance. To tackle those weak points, automatic prompt engineering (APE) is designed for easier prompt generation (Zhou et al., 2022a; Pryzant et al., 2023a; Cheng et al., 2023). Although APE models can automatically enhance the prompt, they have to use massive amounts of human-labeled data (Cheng et al., 2023; Ouyang et al., 2022; Rafailov et al., 2024). Additionally, previous methods failed to construct a flexible, user-friendly and effective APE model (Zhou et al., 2022b; Pryzant et al., 2023b). They face the following three key challenges:
@@
  • New Method. We propose a new method that leverages diversity and quality criteria for prompt data selection. By curating a prompt dataset, we integrate few-shot learning with quality control to construct complementary prompt data. Our method, which employs this curated complementary prompt data to fine-tune LLMs, facilitates effective, efficient, and flexible prompt enhancement.
  • SoTA Performance. (1) High Efficiency. Our PAS model requires only 9000 pairs of prompt complementary data to fine-tune LLMs, making it extremely data-efficient. Additionally, our data generation process is entirely automatic and requires no human labor. Moreover, our PAS model can be integrated into any LLM and solve all tasks, achieving SoTA performance. Therefore, we only need to train one LLM, reducing computational costs. (2) High Flexibility. Our PAS model can be plugged into any LLM, and is model and task agnostic. It achieves SoTA performance across all models and tasks, demonstrating its exceptional flexibility. (3) SoTA Performance on Multiple Benchmarks. Our PAS model achieves SoTA performance across multiple models and comprehensive benchmarks. It also outperforms the previous SoTA model, BPO, under identical experimental settings. Furthermore, it demonstrates superior performance on human evaluation metrics, as shown in Figure 1(b), highlighting the outstanding capabilities and potential applications of the PAS model. Figure 2 further illustrates that our model has significant real-world application potential.
@@ 2.2. Plug-and-Play Systems
In image processing research, plug-and-play systems are commonly applied for its outstanding flexibility. Image reconstruction, denoising, deblurring, image enhancement, and super-resolution are all fields where plug-and-play systems are highly needed. By integrating various image processing modules into a unified framework, plug-and-play systems can flexibly combine different methods to achieve better image processing results. Moreover, this system allows for the easy addition or replacement of new processing modules without redesigning the entire algorithm, significantly improving the efficiency and effectiveness of image processing.
@@ 2.3. Data Quality and Data Selection
: High-quality data can significantly enhance the performance of models (meta llama, 2024). As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation (Bai et al., 2024). Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
@@
: LLMs-based methods were commonly used in data selection (Bai et al., 2024). For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023b) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
@@ 3.1. Prompts Data Collection
3.2. Prompts Complementary Dataset
@@ 4. Experiments
In this section, we first introduce the experimental setups. We then aim to answer the following questions to verify the effectiveness, efficiency, and robustness of our proposed PAS: Q1: Can our PAS achieve SoTA performance compared to previous SoTA methods? Q2: Can our PAS outperform the previous SoTA model with the same base model? Q3: How efficient and flexible is our model compared to previous APE models? Q4: Can PAS achieve SoTA performance in human evaluation, making it a user friendly system? Q5: Can we visualize the advantages of our method? Q6: Do we need data quality selection and regenerate module in our data generation pipeline?
@@ 4.1. Experiments Setting
  • Alpaca-Eval 2.0 (LC): Alpaca-Eval 2.0 LC is a length-controlled version of AlpacaEval designed to mitigate biases related to response length in language model evaluations. By implementing length control, it reduces sensitivity to response length variations, enhancing robustness and interpretability of results. This improvement increases AlpacaEval’s correlation with human judgments, as shown by its higher correlation with Chatbot Arena evaluations.
@@ 4.4. Efficiency and Flexibility of PAS
To address Q3, we compare the data efficiency and flexibility of the PAS approach with SoTA APE models. This comparison underscores PAS’s effectiveness in terms of both data utilization and adaptability.
@@ 4.4.1. Data Usage Comparison
Figure 7 shows the data consumption of PAS compared to other alignment methods, including PPO, DPO, and BPO. The data consumption for each method is as follows:
  • PAS: 9,000
  • BPO: 14,000
  • PPO: 77,000
  • DPO: 170,000
As shown in figure 7, PAS significantly reduces data consumption compared to other methods, demonstrating its efficiency in data usage. It is important to note that ProTeGi and OPRO are not task-agnostic, so they are not included in the data efficiency comparison.
@@ 4.5. Human Evaluation
To address Q4, we conducted a comprehensive evaluation using human evaluators to assess the online performance of our PAS compared to the baseline model without any prompt augmentation. We compared the good same bad (GSB), availability proportion, full mark proportion, and average score.
@@ 4.6. Case Study
Our PAS approach significantly improves this prompt by providing a complementary hint to remind the model that there is a logic trap. With the assistance of PAS, the new response avoids the trap and showcases a multi-step logical reasoning process. This refined answer is clear, accurate, and guides the user through the reasoning process, explaining why only one bird would be on the ground.
@@
Case Study 3
To summarize, the benefits of our PAS method are as follows:
  • Comprehensive and Clear Responses: PAS promotes detailed explanations, ensuring that the response is not only correct but also easy to understand and logically sound, as illustrated in Case Study 3 where PAS provided a thorough analysis of blood pressure changes.